Overall odds of winning a prize in mega-millions is approximately 1/15, and the numbers in each ticket is independent of the others. Richard bought two mega-million tickets, his probability of winning the first OR the second ticket is:

a. 0.0667

b. 0.1333

c. 0.1289
Answer #1

P(winning on a ticket) = 1/15

P(not winning on a ticket) = 1 - 1/15 = 14/15

P(winning the first or the second ticket) = 1 - P(not winning on first and second ticket)

= 1 - (14/15 x 14/15)

= 0.1289

Ans: c. 0.1289